a counter has 4 defective and 3 non defective items.If a sample of 2 items are down one after another without replacement what is the probablity that,the sample will most one defective


1
Expert's answer
2021-08-16T12:19:30-0400

The probability that both items are defective:

"{p_1} = \frac{4}{7} \cdot \frac{3}{6} = \frac{2}{7}"

Probability that only one item is defective:

"{p_2} = \frac{4}{7} \cdot \frac{3}{6} + \frac{3}{7} \cdot \frac{4}{6} = \frac{4}{7}"

Then the wanted probability is

"p(x \ge 1) = {p_1} + {p_2} = \frac{{2 + 4}}{7} = \frac{6}{7}"

Answer: "\frac{6}{7}"